Tropomyosin, a 70-kDa70-kDa muscle protein, is a two‑stranded α‑helical coiled coil. Estimate the length of the molecule.
Answer #1

The molecular weight of two stranded alpha helical coiled coil = 70kDa.

Molecular weight of one alpha helix= 35kDa

Rise per residue of alpha helix = 1.5 Angstrom

Mean molecular weight= 110Da

Now, The residues in each strand = 35 * 1000/ 110 = 318

So, length of the molecule = 318*1.5Angstrom = 477 Angstrom